OpenStudy (anonymous):

Kala drove 200 miles using 9 gallons of gas. At this rate, how many gallons of gas would she need to drive 420 miles? just answer plz

OpenStudy (texaschic101):

200/9 = 420/x (200 to 9 gal = 420 to x gal) cross multiply (200)(x) = (420)(9) 200x = 3780 x = 3780/200 x = 18.9 gallons

OpenStudy (anonymous):

ok thx

OpenStudy (texaschic101):

no problem :)